# Random walks on edge colored random graphs

### v2

@@ 5 The oblivious model: Proof of Theorem 3
Divide $H$ into $\ell=n^{\nu}$ consecutive intervals $I_{1},I_{2},\ldots,I_{\ell}$ of length $L=n^{1-\nu}$, where $30<\nu\leq\varepsilon/$ 3. Fix one such interval $I$. Let $S_{0}$ be the locally tree-like vertices in $I$ and then choose $S$ by greedily choosing a subset of $S_{0}$ where each pair of vertices are at distance $12\sigma+$ 1 apart. Note that $|S|\geq n^{1-\nu-o(1)}$.
@@
$=$ $n^{2\delta}n^{O(\beta/n)}=(\mathbb{E}S_{1})^{2}\left(1+O\left(\frac{1}{n}\right)\right),$
@@
[…]
@@
We require $n^{\theta+(1+\theta)/2+o(1)}\ll L=n^{1-\varepsilon/3}$, which is satisfied by $\theta=(1-\varepsilon)/4$. With this value of $\theta$, w.h.p. each interval $I_{j}$ $j=1,...,\ell$ will contain a vertex unvisited during the free phases. This completes the proof of Theorem 1(b).

### v3

@@ 5 The oblivious model: Proof of Theorem 3
Divide $H$ into $\ell=n^{\nu}$ consecutive intervals $I_{1},I_{2},\ldots,I_{\ell}$ of length $L=n^{1-\nu}$, where $0<\nu\leq\varepsilon/3$. Fix one such interval $I$. Let $S_{0}$ be the locally tree-like vertices in $I$ and then choose $S$ by greedily choosing a subset of $S_{0}$ where each pair of vertices are at distance $2\sigma+1$ apart. Note that $|S|\geq n^{1-\nu-o(1)}$.
@@
$=$ $n^{2\delta}n^{O(\beta/n)}=(\mathbb{E}|S_{1}|)^{2}\left(1+O\left(\frac{1}{n}\right)\right),$
@@
We conclude that w.h.p. each segment $I$ has at least one unvisited vertex in $S_{1}(I)$ when the budget runs out at $t_{1}$. Thus we have to effectively walk around the cycle to reach them all. This requires $\sim n^{2}/2$ time w.h.p. Let $C_{\sc OBV}(\varepsilon)$ be the cover time of an oblivious walk with budget at most $t_{1}=(1-\varepsilon)C(G)$, then
$C_{\sc OBV}(\varepsilon)\approx C(C_{n})\sim\frac{n^{2}}{2}.$
@@
[…]